# A robust parameterized enhanced shift-splitting preconditioner for three-by-three block saddle point problems

### v2

@@ 6 Numerical experiments
  • For $l=32,$ the computed condition numbers of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A},$ and ${\mathscr{P}}_{IBD}^{-1}\mathcal{A}$ are $5.4289e+04,$ $9.5567e+09,$ and $9.5124e+09,$ respectively, which are very large. Whereas Figure 3 illustrate a decreasing trend in the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ (for instance, when $s=50$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=3.4221$) with increasing values of $s.$ This observation highlights that ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ emerges as well-conditioned, consequently the solution obtained by the proposed PESS PGMRES process is more reliable and robust.
@@
The above discussions affirm that the proposed PESS and LPESS preconditioner is efficient, robust, and better well-conditioned with respect to the baseline preconditioners.
@@
Figure 4: Relationship of norm error of solution with increasing noise percentage, employing proposed ${\mathscr{P}}_{PESS}$ for Case II with $s=12$ for $l=16,32,48,64$ and $80$ of Example 1.
Sensitivity analysis of the solution by employing the proposed PESS preconditioner
@@
To study the sensitivity of the solution obtained by employing the proposed PESS PGMRES process to small perturbations on the input data matrices. In the following, we consider the perturbed counterpart $(\mathcal{A}+\Delta\mathcal{A})\tilde{\textbf{u}}=\textbf{d}$ of the three-by-three block SPP (1.1), where the perturbation matrix $\Delta\mathcal{A}$ has same structure as of $\mathcal{A}.$ We construct the perturbation matrix $\Delta\mathcal{A}$ by adding noise to the block matrices $B$ and $C$ of $\mathcal{A}$ as follows:
@@
where $N_{P}$ is the noise percentage and $std(B)$ and $std(C)$ are the standard deviation of $B$ and $C,$ respectively. Here, $randn(m,n)$ is the is an $m$-by-$n$ normally distributed random matrix, generated by the MATLAB command $randn$. We perform the numerical test for $l=16,32,48,64$ and $80$ for Example 1 using the PESS PGMRES process (Case II, $s=12$). The calculated norm error $\|\tilde{\textbf{u}}-\textbf{u}\|_{2}$ among the solution of the perturbed system and the original system with increasing $N_{P}$ from $5\%$ to $40\%$ with step size $5\%$ are displayed in the figure 4. We noticed that with growing $N_{P},$ the norm error of the solution remains consistently less than $10^{-8},$ which demonstrates the robustness of the proposed PESS preconditioner and the solution u is not sensitive to small perturbation on $\mathcal{A}$.
@@
Table 3: Numerical results of GMRES, BD, IBD, SS, RSS, EGSS, PESS and LPESS PGMRES processes for Example 2.
Process h $1/8$ $1/16$ $1/32$ $1/64$
size($\mathcal{A}$) $288$ $1088$ $4224$ $16640$
GMRES IT $158$ $548$ $2048$bf{--}- -
CPU $0.2325$ $0.6960$ $124.7595$ $4545.5072$
RES $078.0113e-$ 07 $078.8023e-$ 07 $079.0887e-$ 07 $064.0481e-$ 06
IBD IT $34$ $35$ $25$ $23$
CPU $\bf 0.0147$ $\bf 0.1166$ $1.9070$ $32.6213$
RES $079.4111e-$ 07 $076.9932e-$ 07 $079.2544e-$ 07 $079.3340e-$ 07
Case I
SS IT $5$ $6$ $9$ $11$
CPU $0.3047$ $0.4422$ $2.4279$ $51.5825$
RES $074.5523e-$ 07 $077.9331e-$ 07 $072.5307e-$ 07 $074.5327e-$ 07
RSS IT $\bf 4$ $\bf 4$ $5$ $5$
CPU $0.2044$ $0.3436$ $1.6047$ $27.6128$
RES $083.4853e-$ 08 $079.2155e-$ 07 $081.7469e-$ 08 $072.7261e-$ 07
EGSS IT $7$ $9$ $11$ $15$
CPU $0.2389$ $0.4051$ $4.4353$ $52.7804$
RES $071.6861e-$ 07 $084.0517e-$ 08 $071.3223e-$ 07 $071.774e-$ 07
PESS IT $4$ $4$ $5$ $5$
CPU $0.2068$ $0.3480$ $1.7633$ $25.2920$
$s=30$ RES $082.3499e-$ 08 $074.2761e-$ 07 $082.1247e-$ 08 $072.0987e-$ 07
LPESS IT $\bf 3$ $\bf{3}$ 4 3
CPU $\bf 0.1990$ $\bf 0.2578$ $\bf 1.4778$ $\bf 17.9730$
$s=30$ RES $071.6606e-$ 07 $078.3803e-$ 07 $104.9289e-$ 10 $075.2658e-$ 07
Case II
SS IT $29$ $38$ $60$ $52$
CPU $0.2999$ $1.1144$ $13.4580$ $213.2381$
RES $077.3560e-$ 07 $079.5790e-$ 07 $079.2515e-$ 07 $079.6425e-$ 07
RSS IT $12$ $13$ $14$ $12$
CPU $0.3797$ $0.6019$ $3.4491$ $50.3476$
RES $073.2456e-$ 07 $075.3086e-$ 07 $076.6600e-$ 07 $074.2834e-$ 07
EGSS IT $9$ $9$ $9$ $7$
CPU $0.2459$ $0.4180$ $2.4320$ $27.9359$
RES $097.9001e-$ 09 $098.6623e-$ 09 $097.4644e-$ 09 $071.2457e-$ 07
PESS IT $5$ $5$ $5$ $5$
CPU $0.2091$ $0.4184$ $1.8118$ $25.8048$
$s=26$ RES $082.3801e-$ 08 $081.1965e-$ 08 $098.4923e-$ 09 $085.9356e-$ 08
LPESS IT $4$ $5$ $\bf 5$ $\bf 5$
CPU $0.2070$ $0.3662$ $\bf 1.5330$ $\bf 23.5997$
$s=26$ RES $072.4250e-$ 07 $128.6182e-$ 12 $115.3930e-$ 11 $108.1253e-$ 10
[…]
@@
Figure 5: Convergence curves for IT versus RES of the PGMRES processes by employing IBD, SS, RSS, EGSS, PESS and LPESS preconditioners in Case II for Example 2.
[…]
Table 3 shows that the GMRES process has a very slow convergence speed and also does not converge for $\textbf{h}=1/64$ within $7000$ iterations. The proposed preconditioners require almost five times fewer iterations compared to the IBD preconditioner for convergence. Moreover, in both cases, the PESS and LPESS preconditioners outperform the SS, RSS, and EGSS preconditioners in terms of IT and CPU times. Notably, IT for the PESS preconditioner, in both cases, remains constant, whereas for SS and EGSS, IT increases as the size of $\mathcal{A}$ increases. Furthermore, in Case II with $l=1/64$, our proposed LPESS preconditioner is approximately $28\%$, $90\%$, $53\%$, and $16\%$ more efficient than the existing IBD, SS, RSS, and EGSS preconditioner, respectively. Similar patterns are observed for other values of $l.$
@@
(c) ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$
(d) ${\mathscr{P}}_{RSS}^{-1}\mathcal{A}$
(e) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
(f) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=26$
(g) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=26$
Figure 6: Spectral distribution of $\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf h}=1/8.$
The convergence curves in Figure 5 demonstrate the rapid convergence of the proposed PESS and LPESS PGMRES processes compared to the baseline IBD, SS, RSS, and EGSS ones in terms of RES versus IT counts. Moreover, Figure 6 illustrates the spectral distributions of the original matrix $\mathcal{A},$ and the preconditioned matrices ${\mathscr{P}}_{IBD}^{-1}\mathcal{A},{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{RSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf{h}}=1/8.$ According to Figure 6, eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ demonstrate a superior clustering compared to the other preconditioned matrices, which leads to a favorable convergence speed for the proposed PESS and LPESS PGMRES processes.

### v3

@@ 7 Numerical experiments
  • For $l=32,$ the computed condition numbers of $\mathcal{A},{\mathscr{P}}_{BD}^{-1}\mathcal{A}$ ${\mathscr{P}}_{IBD}^{-1}\mathcal{A},$ ${{\mathscr{P}}_{MAPSS}^{-1}\mathcal{A}}$ and ${\mathscr{P}}_{SL}^{-1}\mathcal{A}$ are $5.4289e+04,$ $9.5567e+09,$ $9.5124e+09,$ $057.2548e+$ 05 and $4.2852e+09,$ respectively, which are very large. Whereas Figure 4 illustrates a decreasing trend in the condition number of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ (for instance, when $s=50$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=3.4221$) with increasing values of $s.$ This observation highlights that ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ emerges as well-conditioned, consequently, the solution obtained by the proposed PESS PGMRES process is more reliable and robust.
@@
The above discussions affirm that the proposed PESS and LPESS preconditioner is efficient, robust and better well-conditioned with respect to the baseline preconditioners.
@@
Figure 5: Relationship of norm error of solution with increasing noise percentage, employing proposed ${\mathscr{P}}_{PESS}$ for Case II with $s=12$ for $l=16,32,48,64$ and $80$ for Example 1.
@@
Sensitivity analysis of the solution by employing the proposed PESS preconditioner: To study the sensitivity of the solution obtained by employing the proposed PESS PGMRES process to small perturbations on the input data matrices. In the following, we consider the perturbed counterpart $(\mathcal{A}+\Delta\mathcal{A})\tilde{\textbf{u}}=\textbf{d}$ of the three-by-three block SPP (1.1), where the perturbation matrix $\Delta\mathcal{A}$ has same structure as of $\mathcal{A}.$ We construct the perturbation matrix $\Delta\mathcal{A}$ by adding noise to the block matrices $B$ and $C$ of $\mathcal{A}$ as follows:
@@
where $N_{P}$ is the noise percentage and $std(B)$ and $std(C)$ are the standard deviation of $B$ and $C,$ respectively. Here, $randn(m,n)$ is an $m$-by-$n$ normally distributed random matrix generated by the MATLAB command $randn$. We perform the numerical test for $l=16,32,48,64$ and $80$ for Example 1 using the PESS PGMRES process (Case II, $s=12$). The calculated norm error $\|\tilde{\textbf{u}}-\textbf{u}\|_{2}$ among the solution of the perturbed system and the original system with increasing $N_{P}$ from $5\%$ to $40\%$ with step size $5\%$ are displayed in the Figure 5. We noticed that with growing $N_{P},$ the norm error of the solution remains consistently less than $10^{-8},$ which demonstrates the robustness of the proposed PESS preconditioner and the solution u is insensitive to small perturbation on $\mathcal{A}$.
@@
Table 5: Numerical results of GMRES, IBD, MAPSS, SL, SS, RSS, EGSS, RPGSS, PESS and LPESS PGMRES processes for Example 2.
Process h $1/8$ $1/16$ $1/32$ $1/64$ $1/128$
size($\mathcal{A}$) $288$ $1088$ $4224$ $16640$ $66048$
[…]
CPU $0.2325$ $0.6960$ $124.7595$ $4545.5072$ $9396.6958$
RES $078.0113e-$ 07 $078.8023e-$ 07 $079.0887e-$ 07 $064.0481e-$ 06 $068.3021e-$ 06
IBD IT $34$ $35$ $25$ $23$ $21$
CPU $\bf 0.0147$ $\bf 0.1166$ $1.9070$ $32.6213$ $1664.6991$
RES $079.4111e-$ 07 $076.9932e-$ 07 $079.2544e-$ 07 $079.3340e-$ 07 $077.3550e-$ 07
MAPSS IT $11$ $7$ $4$ $4$ $4$
CPU $0.2061$ $0.3453$ $1.8717$ $26.6600$ $482.0390$
RES $085.2642e-$ 08 $071.1366e-$ 07 $087.1116e-$ 08 $078.2780e-$ 07 $092.1189e-$ 09
SL IT $6$ $6$ $5$ $5$ $4$
CPU $0.2130$ $0.2929$ $1.8697$ $27.1124$ $516.3012$
RES $076.6518e-$ 07 $081.1953e-$ 08 $089.5638e-$ 08 $112.1528e-$ 11 $102.9479e-$ 10
Case I
SS IT $5$ $6$ $9$ $11$ $11$
CPU $0.3047$ $0.4422$ $2.4279$ $51.5825$ $1245.2019$
RES $074.5523e-$ 07 $077.9331e-$ 07 $072.5307e-$ 07 $074.5327e-$ 07 $079.8497e-$ 07
RSS IT $\bf 4$ $\bf 4$ $5$ $5$ $6$
CPU $0.2044$ $0.3436$ $1.6047$ $27.6128$ $720.4234$
RES $083.4853e-$ 08 $079.2155e-$ 07 $081.7469e-$ 08 $072.7261e-$ 07 $093.2621e-$ 09
EGSS IT $7$ $9$ $11$ $15$ $5$
CPU $0.2389$ $0.4051$ $4.4353$ $52.7804$ $744.5983$
RES $071.6861e-$ 07 $084.0517e-$ 08 $071.3223e-$ 07 $071.774e-$ 07 $079.1207e-$ 07
RPGSS IT $4$ $4$ $4$ $5$ $5$
CPU $0.2105$ $0.3883$ $2.2981$ $28.1748$ $544.8828$
RES $092.8700e-$ 09 $085.1131e-$ 08 $073.3646e-$ 07 $098.8663e-$ 09 $071.4981e-$ 07
PESS IT $4$ $4$ $5$ $5$ $4$
CPU $0.2068$ $0.3480$ $1.7633$ $25.2920$ $484.2796$
$s=30$ RES $082.3499e-$ 08 $074.2761e-$ 07 $082.1247e-$ 08 $072.0987e-$ 07 $074.3859e-$ 07
LPESS IT $\bf 3$ $\bf{3}$ 4 3 44{4}4
CPU $\bf 0.1990$ $\bf 0.2578$ $\bf 1.4778$ $\bf 17.9730$ $457.7446$
$s=30$ RES $071.6606e-$ 07 $078.3803e-$ 07 $104.9289e-$ 10 $075.2658e-$ 07 $083.3191e-$ 08
Case II
SS IT $29$ $38$ $60$ $52$ $55$
CPU $0.2999$ $1.1144$ $13.4580$ $213.2381$ $4980.7387$
RES $077.3560e-$ 07 $079.5790e-$ 07 $079.2515e-$ 07 $079.6425e-$ 07 $079.6948e-$ 07
RSS IT $12$ $13$ $14$ $12$ $12$
CPU $0.3797$ $0.6019$ $3.4491$ $50.3476$ $1361.8902$
RES $073.2456e-$ 07 $075.3086e-$ 07 $076.6600e-$ 07 $074.2834e-$ 07 $071.2733e-$ 07
EGSS IT $9$ $9$ $9$ $7$ $4$
CPU $0.2459$ $0.4180$ $2.4320$ $27.9359$ $419.1294$
RES $097.9001e-$ 09 $098.6623e-$ 09 $097.4644e-$ 09 $071.2457e-$ 07 $087.6469e-$ 08
RPGSS IT $6$ $7$ $7$ $5$ $4$
CPU $0.2154$ $0.4096$ $1.8893$ $23.2802$ $423.0121$
RES $085.2642e-$ 08 $071.1366e-$ 07 $087.1116e-$ 08 $078.2780e-$ 07 $092.1189e-$ 09
PESS IT $5$ $5$ $5$ $5$ $\bf 3$
CPU $0.2091$ $0.4184$ $1.8118$ $25.8048$ $\bf 365.3128$
$s=26$ RES $082.3801e-$ 08 $081.1965e-$ 08 $098.4923e-$ 09 $085.9356e-$ 08 $074.2039e-$ 07
LPESS IT $4$ $5$ $\bf 5$ $\bf 5$ $\bf 3$
CPU $0.2070$ $0.3662$ $\bf 1.5330$ $\bf 23.5997$ $\bf 359.1330$
$s=26$ RES $072.4250e-$ 07 $128.6182e-$ 12 $115.3930e-$ 11 $108.1253e-$ 10 $075.8345e-$ 07
[…]
@@
Table 6: Numerical results of PESS-I, LPESS-I, PESS-II and LPESS-II PGMRES processes for Example 2.
Process h $1/8$ $1/16$ $1/32$ $1/64$ $1/128$
size($\mathcal{A}$) $288$ $1088$ $4224$ $16640$ $66048$
PESS-I IT $5$ $5$ $6$ $6$ $5$
($s=1,$ $\Lambda_{1}=0.001I,$ CPU $0.30867$ $0.46688$ $2.48254$ $46.52887$ $577.9420$
$\Lambda_{2}=0.1I,$ $\Lambda_{3}=0.001I$) RES $081.4125e-$ 08 $071.3379e-$ 07 $083.8821e-$ 08 $072.6061e-$ 07 $079.7383e-$ 07
LPESS-I IT $4$ $4$ $4$ $5$ $3$
($s=1,$ $\Lambda_{2}=0.1I,$ CPU $0.1829$ $0.3319$ $1.2774$ $19.1090$ $349.6728$
$\Lambda_{3}=0.001I$) RES $093.5507e-$ 09 $083.4257e-$ 08 $077.0375e-$ 07 $092.3073e-$ 09 $077.1453e-$ 07
PESS-II IT $3$ $4$ $4$ $4$ $4$
($s=s_{est},$ $\Lambda_{1}=A,$ CPU $0.18470$ $0.39536$ $1.32148$ $17.22840$ $485.01767$
$\Lambda_{2}=\beta_{est}I,$ $\Lambda_{3}=10^{-4}CC^{T}$) RES $079.6100e-$ 07 $075.6520e-$ 07 $087.8750e-$ 08 $083.9903e-$ 08 $096.9694e-$ 09
LPESS-II IT $6$ $6$ $5$ $5$ $3$
($s=s_{est},$ $\Lambda_{2}=\beta_{est}I,$ CPU $0.1656$ $0.3777$ $1.4341$ $19.0658$ $377.2001$
$\Lambda_{3}=10^{-4}CC^{T}$) RES $072.5227e-$ 07 $075.1622e-$ 07 $084.8924e-$ 08 $083.2751e-$ 08 $077.9950e-$ 07
To generate the matrices $A\in{\mathbb{R}}^{n\times n}$ and $B\in{\mathbb{R}}^{m\times n}$ of the system (1.1), the discretization task of the Stokes equation (7.1) is accomplished by the IFISS software developed by Elman et al. [3]. Following [3] with grid parameters $\textbf{h}=1/8,1/16,1/32,1/64,{1/128}$, we get $n=2(1+1/\bf{h})^{2}$ and $m=1/\textbf{h}^{2}.$ To make the system of equations in (1.1) not too ill-conditioned and not too sparse, we construct the block matrix $C=\left[\begin{matrix}\Pi&randn(p,m-p)\end{matrix}\right]$, where $\Pi=\mathrm{diag}(1,3,5,\ldots,2p-1)$ and $p=m-2.$ Additionally, to ensure the positive definiteness of the matrix $A$, we add $0.001I$ with $A.$
@@
(e) $\textbf{h}=1/128$
Figure 6: Convergence curves for IT versus RES of the PGMRES processes by employing IBD, MAPSS, SL, SS, RSS, EGSS, RPGSS, PESS and LPESS preconditioners in Case II for Example 2.
@@
(c) ${\mathscr{P}}_{MAPSS}^{-1}\mathcal{A}$
(d) ${\mathscr{P}}_{SL}^{-1}\mathcal{A}$
(e) ${\mathscr{P}}_{SS}^{-1}\mathcal{A}$
(f) ${\mathscr{P}}_{RSS}^{-1}\mathcal{A}$
(g) ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A}$
(h) ${\mathscr{P}}_{RPGSS}^{-1}\mathcal{A}$
(i) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=26$
(j) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=26$
Figure 7: Spectral distributions of $\mathcal{A},{\mathscr{P}}_{IBD}^{-1}\mathcal{A},{{\mathscr{P}}_{MAPSS}^{-1}\mathcal{A},{\mathscr{P}}_{SL}^{-1}\mathcal{A},}{\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{EGSS}^{-1}\mathcal{A},{{\mathscr{P}}_{RPGSS}^{-1}\mathcal{A},}{\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf h}=1/8.$
[…]
Parameter selection: Parameter choices for the proposed preconditioners are made in two cases as in Example 1. However, in Case I, we take $\alpha=0.01,\beta=0.1,$ $\Lambda_{1}=0.01I$ and $\Lambda_{2}=0.1I.$ The parameter $\alpha$ in Case I is taken as in [36] and the parameters in Case II are taken as in [39]. For the IBD preconditioner, the matrices $\widehat{A}$ and $\widehat{S}$ and for MAPSS preconditioner $\alpha$ and $\beta$ are constructed as in Example 1.
[…]
(a) ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ with $s=26$
(b) ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ with $s=26$
Figure 8: Spectral bounds for ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf h}=1/8$ for Example 2.
Results using parameters selection strategy in Section 6: To demonstrate the effectiveness of the proposed PESS and LPESS preconditioners using the parameters discussed in Section 6, we present the numerical test results for PESS-I, LPESS-I, PESS-II and LPESS-II in Table 6 as in Example 1. Comparing the results in Tables 5 and 6, we observe that the parameter selection strategy described in Section 6 is effective.
Table 7: Spectral bounds for ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for case II with $l=16$ for Example 2
Real eigenvalue $\lambda$ Non-real eigenvalue $\lambda$
Bounds of Theorem 4.2 Bounds of Theorem 4.3
${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ contained in $0.0334\leq|\lambda|\leq 0.375$
the interval $3.3624\times 10^{-6}\leq\Re(\lambda/(1-s\lambda))\leq 0.5$
$(0,0.0370]$ $|\Im(\lambda/(1-s\lambda))|\leq 31.6386$
Bounds of Theorem 5.1 Bounds of Theorem 5.1
${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ contained in $0.0030\leq|\lambda|\leq 0.0385$
the interval $4.6722\times 10^{-5}\leq|\lambda-\frac{1}{s}|\leq 0.0355$
$(0.00552,0.0270]$
Convergence curves: The convergence curves in Figure 6 demonstrate the rapid convergence of the proposed PESS and LPESS PGMRES processes compared to the IBD, MAPSS, SL, SS, RSS, EGSS and RPGSS in terms of RES versus IT counts.
Spectral distributions: Figure 7 illustrates the spectral distributions of the original matrix $\mathcal{A},$ and the preconditioned matrices ${\mathscr{P}}_{IBD}^{-1}\mathcal{A},$ ${{\mathscr{P}}_{MAPSS}^{-1}\mathcal{A},}$ ${{\mathscr{P}}_{SL}^{-1}\mathcal{A},}$ ${\mathscr{P}}_{SS}^{-1}\mathcal{A},{\mathscr{P}}_{RSS}^{-1}\mathcal{A},$ ${\mathscr{P}}_{EGSS}^{-1}\mathcal{A},$ ${{\mathscr{P}}_{PRGSS}^{-1}\mathcal{A},}$ ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ for Case II with ${\bf{h}}=1/8.$ According to Figure 7, eigenvalues of ${\mathscr{P}}_{PESS}^{-1}\mathcal{A}$ and ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ demonstrate a superior clustering compared to the other preconditioned matrices, which leads to a favorable convergence speed for the proposed PESS and LPESS PGMRES processes.
Spectral bounds: In Figure 8, we draw the spectral bounds of Theorems 4.1, 4.2 and 5.1 for PESS and LPESS preconditioned matrices. In Figure 8(a), $|\lambda-1|=1$ of Theorem 4.1 is drawn by the green unit circle, while the points in blue indicate the bounds from Theorem 4.2. Moreover, we draw the circles $C_{1}$ (in red), $C_{2}$ (in blue), $C_{3}$ (in black) and $C_{4}$ (in green) in Figure 8(b). We can observe that the eigenvalues of the preconditioned matrix ${\mathscr{P}}_{LPESS}^{-1}\mathcal{A}$ lie in the intersection of the annulus of the circles $C_{1}$ and $C_{2},$ and the annulus of the circles $C_{3}$ and $C_{4}.$ Additionally, the eigenvalue bounds derived in Theorems 4.2, 4.3 and 5.1 are listed in Table 7.
@@
Condition number analysis: Furthermore, the system (1.1) exhibits ill-conditioning nature with $\kappa(\mathcal{A})=5.0701e+05.$ While for Case I with $\mathbf{h}=1/16$, $\kappa({\mathscr{P}}_{PESS}^{-1}\mathcal{A})=1.3353$ and $\kappa({\mathscr{P}}_{LPESS}^{-1}\mathcal{A})=1.2056,$ indicating that the proposed PESS and LPESS preconditioned systems are well-conditioned, ensuring an efficient and robust solution.